# QA-MDT: Quality-aware Masked Diffusion Transformer for Enhanced Music Generation

### v2

@@ 4 Method
The issue can stem from the original data being improperly labeled with terms such as ’speech, car’ from datasets like AudioSet (Gemmeke et al. 2017) and also may be because of desperately missing of the original labels.
@@ 5 Experimental Setup
Datasets
Train datasets.
We used the following databases for our training: AudioSet Music Subset (ASM) (Gemmeke et al. 2017), MagnaTagTune (MTT) (Law et al. 2009), Million Song Dataset (MSD) (Bertin-Mahieux et al. 2011), Free Music Archive (FMA) (Defferrard et al. 2016), and an additional datasetWe use 55k music tracks from https://pixabay.com, which is a large scale copyright free dataset.. Each track in these databases was clipped to 10-second segments and sampled at 16kHz to ensure uniformity across the dataset. The final training set was developed through a process of caption refinement, as detailed in Section 4. Finally, we got our training set totaling 12.5k hours of diverse music data. The specific composition of these datasets is further elaborated in the Appendix.
Evaluation datasets.
For comparison with prior work, we evaluate our model on the widely used MusicCaps benchmark (Agostinelli et al. 2023) and the Song-Describer-Dataset (Manco et al. 2023). MusicCaps consists of 5.5K $10.24$-second clips that include high-quality music descriptions from ten musicians, and the Song-Describer-Dataset is made up of 706 licensed high quality music recordings.
Models and Hyperparameters
@@
Each 10.24-second audio clip, sampled at 16 kHz, is initially transformed into a $64\times 1024$ mel-spectrogram with mel-bins of 64, hop-length of 160 and window length of 1024. Subsequently, this spectrogram is compressed into a $16\times 128$ latent representation $\mathcal{M}_{spec}$ using a Variational Autoencoder (VAE) pretrained with AudioLDM 2 (Liu et al. 2023b) with series of quantization loss and adversarial loss. This latent representation is utilized for adding noise and training purposes. Finally, we employ a pretrained Hifi-GAN (Kong, Kim, and Bae 2020) to reconstruct the waveform from the generated mel-spectrogram.
@@
We utilize the LP-MusicCaps (Doh et al. 2023) caption model for ASM, FMA, and subsets of MTT and MSD that have weak or no captions. We use the official checkpoint from LAION-CLAP (Wu* et al. 2023)music_speech_audioset_epoch_15_esc_89.98.pt for text-to-text and text-to-audio similarity calculations. Based on small scale subjective experientment, thresholds are set at $\rho_{1}=\rho_{2}=0.1$ to ensure any generated text or original caption not aligned well with the corresponding waveform is filtered out. Additionally, after filtering, generated text that fall below a threshold of $\rho_{3}=0.25$ are merged with original tags with the prompt: Merge this music caption “generated caption” with the ground truth tags “original tags”. Do not add any imaginary elements and try to keep the modified caption refined and accurate. We use FLAN-T5-large (Peebles and Xie 2023) as text encoder for all models.
@@
We train our diffusion model with three backbones for comparison: U-Net (Ronneberger, Fischer, and Brox 2015) based at 1.0B parameters, code based on AudioLDM (Liu et al. 2023c); and our proposed Quality-aware Masked Diffusion Transformer (QA-MDT) based at 675M parameters with $N=20$ encoder layers and $M=8$ decoder layers. Flash attention (Dao et al. 2022) from the xFormers package (Lefaudeux et al. 2022) is utilized to improve training speed and memory usage for self-attention and cross-attention stages. We study the impact of the patch size and overlap size in the Appendix , and apply a patch size of $1\times 4$ without overlap for the training of our final model. We train on 10-second audio crops sampled at random from the full track, maintaining a total batch size of 64 on four NVIDIA A100 GPUs, with gradient clipping of 1.0, learning rate of 8e-5, 2000 steps of linear warmup without weight decay, and a condition drop of 0.1 during training. The final model was trained for a total of seven days (38.5k steps). During inference, we use Denoising Diffusion Implicit Models (DDIM) (Song, Meng, and Ermon 2020) with 200 steps and a guidance scale of 3.5, consistent with AudioLDM (Liu et al. 2023c). We begin by presenting our approach to refining captioning, which includes the capability for quality awareness, transitioning from text-level control to token-level control. Finally, we compare proposed model with previous works subjectively and objectively.
Evaluation Metrics
We evaluate the proposed method using objective metrics, including the Fréchet Audio Distance (FAD) (Kilgour et al. 2018), Kullback-Leibler Divergence (KL), Inception Score (IS)All above metrics are computed using the audioldm_eval library (Liu et al. 2023c), ensuring standardized evaluation.. We also utilize pseudo-MOS scoring model (Ragano, Benetos, and Hines 2023) to estimate generation quality, with more accurate assessments derived from subjective metrics. For the human studies, we invite human raters to evaluate two aspects of the audio samples (i) overall quality (Ovl), and (ii) relevance to the text input (Rel). For the overall quality test, raters were asked to rate the perceptual quality of the provided samples with 1 to 5. For the text relevance test, raters were asked to rate the match between audio and text on a scale from 1 to 5. Our evaluators consist of people from various backgrounds, including professional music producers, video editors who need to use soundtracks, staff responsible for the operations of music products and complete beginners who are unfamiliar with related knowledge. We evaluate randomly sampled file, where each sample was evaluated by at least 5 raters.
@@ 6 Results
Quality Awareness
Figure 3: (Left) Five VQ-MOS distribution curves are obtained by concurrently using text quality prefixes and quality tokens as controls on the MTT-FS, with quantized MOS levels ranging from 1 to 5 serving as control constraint inferences. The distribution of the training set is normalized by each sample’s duration, colored lines represent thresholds of quantized p-MOS tokens during training. (Right) The effect of using quality text prefixes during training is shown, showcasing testing results on four metrics: FAD, KL, IS, and p-MOS, while gray lines for quantized p-MOS.
@@
This subsection explores the effects and interactions of model control over quality tokens and quality text prefixes during the training phase, as well as their comparative effects across different models. In our previous MTT dataset of 1,000 test pairs, we filtered out pairs labeled with low quality or quality is poor to avoid confusion when applying quality prefixes, resulting in a new subset of 519 entries, which we refer to as the MTT Filter Set (MTT-FS). Figure 4 illustrates the impact of different quality prefixes during inference when quality is used as a text prefix during training for U-Net and MDT-based backbones. It was observed that U-Net, when inferred with different quality prefixes, showed only minor changes in p-MOS scores and did not adhere to the threshold set during training. In contrast, MDT demonstrated better learning of quality information from prefixes, achieving p-MOS scores significantly higher than those of U-Net and the test set. Additionally, by decoupling quality information from the training set, we achieved superior diversity and generalizability compared to training and inference without quality text prefixes. Given that quality tokens are specifically designed for the Transformer architecture, Figure 3 (left) shows the controlled outcomes when different quality tokens are used after integrating quantified quality as a token during training. Remarkably, using quality tokens alone provided more precise and accurate p-MOS score control. In our ablation study, we compared the effects of using only text prefixes against combining both approaches. As shown in Figure 3 (right), as the quantized control level gradually increased, the model steadily improved in p-MOS scores, which represent the quality of generation. Concurrently, FAD and KL also progressively optimized until a turning point at level 4, where a higher average p-MOS was achieved than when solely using prefixes. This turning point may be due to the scarcity of examples with quality level 5 in the dataset. Moreover, by combining two types of quality information injection, the refined decoupling and interaction allowed the model to more accurately perceive audio data quality features during training, leading to significant reductions in FAD and KL compared to using only one of the methods.
We also compare our approach with the traditional “negative prompt” strategy in Appendix, highlighting our approach’s significant improvement in quality and reduction in FAD.
Impact of Music Caption Refinement
@@
We conducted our ablation study on a subset of our training set, which includes ASM and FMA, totaling approximately 3,700 hours and 1.1 million clips. For evaluation, we utilized an out-of-domain set with 1,000 samples randomly selected from MTT (Law et al. 2009). Table  1 compares the model’s performance using different textual representations: sentences formed by merging original tags with commas ($\mathcal{D}_{o}$), generated captions ($\mathcal{D}_{g}$), and generated captions refined through filtering and fusion ($\mathcal{D}_{\text{merge}}$). During the filtering and fusion stage, 8.9% of the generated captions were filtered out, and 15.1% were fused with original tags using ChatGPT. Each model underwent training for 60,000 steps with a batch size of 64.
@@
Compared with Previous Methods
We compared our proposed method with the following representative previous methods: AudioLDM 2 (Liu et al. 2023b), Mousai (Schneider et al. 2023) and Jen-1 (Li et al. 2023) which model music using spectral latent spaces, MusicLM (Agostinelli et al. 2023), and MusicGen (Copet et al. 2024), which focus on modeling discrete representations.
We re-inferred AudioLDM2-Music and MusicGen-1.5B using their official checkpoints to compare additional metrics under the same environment. The results are presented in Table 2. For Ours (U-Net), we inferred all text with the prefix “high quality”, while for Ours (QA-MDT), we used the same prefix along with a p-MOS quality token set to level 5. When calculating the CLAP score, we evaluated the generated music with original prompt, which did not include any quality prefix. The experimental results show significant advantages in both subjective and objective metrics for our models. Since KL divergence measures the distance between audio samples, higher quality audio often results in deviations from the original waveform of Musiccaps, which can lead to lower performance. Although Ours (U-Net) showed a slight FAD advantage on the Song-Describer-Dataset, this may be due to instabilities arising from the small dataset, and we further demonstrated the superiority of QA-MDT in subsequent subjective experiments. Additionally, since MusicGen was trained on non-vocal tracks, it may underperform on captions that include vocals.
Based on subjective evaluation shown in Table 3, our proposed method significantly improves overall audio quality and text alignment, thanks to the label optimization for large music datasets and the quality-aware training strategy. By analyzing the backgrounds of the evaluators and their corresponding results, we can also see that for beginners, the comparison between different systems is not sensitive, which is related to their lack of music background experience and knowledge. However, from the perspective of our method in product operators, video editors, and audio producers, our method offers considerable enhancements, underscoring its potential value to audio industry professionals.
@@ 7 Conclusion and Discussion
In this study, we identify the challenges posed by large-scale uneven audio quality and unaligned textual annotations in the music generation domain, which hinder the development of diffusion-based TTM generation. By employing a novel quality awareness learning approach based on p-MOS, along with masked diffusion transformer as the backbone for the diffusion process, we achieve enhanced generation quality and musicality for music generation.
@@ Appendix A Appendix
Comparing with Using “Negative Prompt”
@@
$D_{\theta}^{\text{CFG}}(x,t,y)=D_{\theta}(x,t,y)+w\left(D_{\theta}(x,t,y)-D_{\theta}(x,t,y_{neg})\right),$ (11)
@@
Based on the TableA.3, we found that any form of quality guidance improves the model’s generative performance. However, previous attempts to improve quality relied on the rare instances of ”low quality” in the dataset. This necessitated the careful design of numerous negative prompts to avoid generating low-quality results. Furthermore, the text embedding of ”low quality” might not be well disentangled during training, leading to an suboptimal results.

### v3

@@ 4.3 Music Caption Refinement
The issue can stem from the original data being improperly labeled with terms such as ’speech, car’ from datasets like AudioSet (Gemmeke et al., 2017) and also may be because of desperately missing of the original labels.
@@ 5 Experimental Setup
5.1 Datasets
For training, we used the following databases for our training: AudioSet Music Subset (ASM) (Gemmeke et al., 2017), MagnaTagTune (MTT) (Law et al., 2009), Million Song Dataset (MSD) (Bertin-Mahieux et al., 2011), Free Music Archive (FMA) (Defferrard et al., 2016), and an additional dataset [^1] We use 55k music tracks from https://pixabay.com, which is a large scale copyright free dataset.. Each track in these databases was clipped to 10-second segments and sampled at 16kHz to ensure uniformity across the dataset. The final training set was developed through a process of caption refinement, as detailed in Section 4.3. Finally, we got our training set totaling 12.5k hours of diverse music data. The specific composition of these datasets is further elaborated in the Appendix. For evaluation, we test our model on the widely used MusicCaps benchmark (Agostinelli et al., 2023) and the Song-Describer-Dataset (Manco et al., 2023). MusicCaps consists of 5.5K 10.24-second clips sourced from YouTube, each accompanied by high-quality music descriptions provided by ten musicians. The Song-Describer Dataset is made up of 706 licensed high quality music recordings.
5.2 Models and Hyperparameters
@@
Each 10.24-second audio clip, sampled at 16 kHz, is initially transformed into a $64\times 1024$ mel-spectrogram with mel-bins of 64, hop-length of 160 and window length of 1024. Subsequently, this spectrogram is compressed into a $16\times 128$ latent representation $\mathcal{M}_{spec}$ using a Variational Autoencoder (VAE) pretrained with AudioLDM 2 (Liu et al., 2023a) with series of quantization loss and adversarial loss. We use pretrained Hifi-GAN (Kong et al., 2020) vocoder to reconstruct the waveform from the generated mel-spectrogram.
@@
We utilize the LP-MusicCaps (Doh et al., 2023) caption model for ASM, FMA, and subsets of MTT and MSD that have weak or no captions. We use the official checkpoint from LAION-CLAP (Wu* et al., 2023)222music_speech_audioset_epoch_15_esc_89.98.pt for text-to-text and text-to-audio similarity calculations. Based on small scale subjective experientment, thresholds are set at $\rho_{1}=\rho_{2}=0.1$ to ensure any generated text or original caption not aligned well with the corresponding waveform is filtered out. Additionally, after filtering, generated text that fall below a threshold of $\rho_{3}=0.25$ are merged with original tags with the prompt: Merge this music caption “generated caption” with the ground truth tags “original tags”, and do not add any imaginary elements.. We use FLAN-T5-large (Peebles and Xie, 2023) as text encoder for all models.
@@
We train our diffusion model with three backbones for comparison: U-Net (Ronneberger et al., 2015) based at 1.0B parameters; our proposed Quality-aware Masked Diffusion Transformer (QA-MDT) with $N=20$ encoder layers and $M=8$ decoder layers at 675M. We study the impact of the patch size and overlap size in the Appendix, and apply a patch size of $1\times 4$ without overlap for the training of our final model. We train on 10.24-second audio crops sampled at random from the full track, maintaining a total batch size of 64, learning rate of 8e-5, and a condition drop of 0.1 during training. The final model was trained for a total of 38.5k steps. During inference, we use Denoising Diffusion Implicit Models (DDIM) (Song et al., 2020) with 200 steps and a guidance scale of 3.5, consistent with AudioLDM (Liu et al., 2023b). We begin by presenting our approach to refining captioning, which includes the capability for quality awareness, transitioning from text-level control to token-level control. Finally, we compare proposed model with previous works subjectively and objectively.
5.3 Evaluation Metrics
We evaluate the proposed method using objective metrics, including the Fréchet Audio Distance (FAD) (Kilgour et al., 2018), Kullback-Leibler Divergence (KL), Inception Score (IS) [^3] We strictly follow the comparison method and evalution code in AudioLDM2 and ensure that the indicators in the paper follow consistent sampling rates and durations for fair comparison. All above metrics are computed using the audioldm_eval library (Liu et al., 2023b), ensuring standardized evaluation.. We also utilize pseudo-MOS scoring model (Ragano et al., 2023) to estimate generation quality, with more accurate assessments derived from subjective metrics.
For the human studies, we invite human raters to evaluate two aspects of the audio samples (i) overall quality (Ovl), and (ii) relevance to the text input (Rel). For the overall quality test, raters were asked to rate the perceptual quality of the provided samples with 1 to 5. For the text relevance test, raters were asked to rate the match between audio and text on a scale from 1 to 5. Our evaluators consist of people from various backgrounds, including professional music producers, video editors, staff responsible for the operations of music products and complete beginners who are unfamiliar with related knowledge. We evaluate randomly sampled file and each sample was evaluated by at least 5 raters.
@@ 6 Results
6.1 Quality Awareness
Figure 3: (Left) Five p-MOS distribution curves are obtained by concurrently using text quality prefixes and quality tokens as controls on the MTT-FS, with quantized MOS levels ranging from 1 to 5 serving as control constraint inferences. The distribution of the training set is normalized by each sample’s duration, colored lines represent thresholds of quantized p-MOS tokens during training. (Right) The effect of using quality text prefixes during training is shown, showcasing testing results on FAD and p-MOS, while gray lines for quantized p-MOS threshold.
@@ 6.1 Quality Awareness
This subsection explores the effects and interactions of model control over quality tokens and quality text prefixes during the training phase, as well as their comparative effects across different models. In our previous MTT dataset of 1,000 test pairs, we filtered out pairs labeled with low quality or quality is poor to avoid confusion when applying quality prefixes, resulting in a new subset of 519 entries, which we refer to as the MTT Filter Set (MTT-FS). Figure 4 illustrates the impact of different quality prefixes during inference when quality is used as a text prefix during training for U-Net and MDT-based backbones. It was observed that U-Net, when inferred with different quality prefixes, showed only minor changes in p-MOS scores and did not adhere to the threshold set during training. In contrast, MDT demonstrated better learning of quality information from prefixes, achieving p-MOS scores significantly higher than those of U-Net and the test set. Additionally, by decoupling quality information from the training set, we achieved better FAD (5.602 vs 5.757) and higher p-MOS (4.039 vs 3.796) compared to training and inference without quality text prefixes. Given that quality tokens are specifically designed for the Transformer architecture, Figure 3 (left) shows the controlled outcomes when different quality tokens are used after integrating quantified quality as a token during training. Remarkably, using quality tokens alone provided more precise and accurate p-MOS score control. In our ablation study, we compared the effects of using only text prefixes against combining both approaches. As shown in Figure 3 (right), as the quantized control level gradually increased, the model steadily improved in p-MOS scores, which represent the quality of generation. Concurrently, FAD and KL also progressively optimized until a turning point at level 4, where a higher average p-MOS was achieved than when solely using prefixes. This turning point may be due to the scarcity of examples with quality level 5 in the dataset. Moreover, by combining two types of quality information injection, the refined decoupling and interaction allowed the model to more accurately perceive audio data quality features during training, leading to significant reductions in FAD and KL compared to using only one of them.
We also compare our approach with the traditional “negative prompt” strategy (Liu et al., 2023a) in Appendix, highlighting our approach’s significant improvement in quality and reduction in FAD.
6.2 Impact of Music Caption Refinement
@@
We conducted our ablation study on a subset of our training set, which includes ASM and FMA, totaling approximately 3,700 hours and 1.1 million clips. For evaluation, we utilized an out-of-domain set with 1,000 samples randomly selected from MTT (Law et al., 2009). Table  1 compares the model’s performance using different textual representations: sentences formed by merging original tags with commas ($\mathcal{D}_{o}$), generated captions ($\mathcal{D}_{g}$), and generated captions refined through filtering and fusion ($\mathcal{D}_{\text{merge}}$). During the filtering and fusion stage, 8.9% of the generated captions were filtered out, and 15.1% were fused with original tags using ChatGPT. Each model underwent training for 60,000 steps with a batch size of 64.
@@
6.3 Compared with Previous Methods
We compared our proposed method with the following representative previous methods: AudioLDM 2 (Liu et al., 2023a), Mousai (Schneider et al., 2023) and Jen-1 (Li et al., 2023) which model music using spectral latent spaces, MusicLM (Agostinelli et al., 2023), and MusicGen Copet et al. (2024), which focus on modeling discrete representations.
We re-inferred AudioLDM2-Music and MusicGen-1.5B using their official checkpoints to compare additional metrics under the same environment. The results are presented in Table 2. For Ours (U-Net), we inferred all text with the prefix “high quality”, while for Ours (QA-MDT), we used the same prefix along with a p-MOS quality token set to level 5. When calculating the CLAP score, we evaluated the generated music with original prompt, which did not include any quality prefix. The experimental results show significant advantages in both subjective and objective metrics for our models. Since KL divergence measures the distance between audio samples, higher quality audio often results in deviations from the original waveform of Musiccaps, which can lead to lower performance. Although Ours (U-Net) showed a slight FAD advantage on the Song-Describer-Dataset, this may be due to instabilities arising from the small scale test dataset, and we further demonstrated the superiority of QA-MDT in subsequent subjective experiments. Additionally, since MusicGen was trained on non-vocal tracks, it may underperform on captions that include vocals.
Based on subjective evaluation shown in Table 3, our proposed method significantly improves overall audio quality and text-audio consistency, thanks to the label optimization for large music datasets and the quality-aware training strategy. By analyzing the backgrounds of the evaluators and their corresponding results, we can also see that for beginners, the comparison between different systems is not sensitive, which is related to their lack of music background experience and knowledge. However, from the perspective of our method in product operators, video editors, and audio producers, our method offers considerable enhancements, underscoring its potential value to audio industry professionals.
@@ 7 Conclusion and discussion
In this study, we address the key challenges in the music generation domain, including model architecture design, large-scale uneven audio quality, and unaligned textual annotations, all of which impede the progress of TTM. By introducing a novel quality-aware training approach, along with MDT backbone and a three-stage text alignment scheme on large-scale open-source music dataset, we improve generation quality, musicality, and text alignment. In the future, we aim to further enhance and expand our model to achieve long-duration, high-sampling-rate, controllable, and highly interactive music generation. Additionally, we are committed to open-sourcing our model and checkpoints.
Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant 62171427.
@@ A.2 Patchify Strategy
A.3 Comparing with Using “Negative Prompt”
@@
$\small D_{\theta}^{\text{CFG}}(x,t,y)=D_{\theta}(x,t,y)+w\left(D_{\theta}(x,t,y)-D_{\theta}(x,t,y_{neg})\right),$ (11)
@@
Based on the TableA.3, we found that any form of quality guidance improves the model’s generative performance. However, previous attempts to improve quality relied on the rare instances of ”low quality” in the dataset. This necessitated the careful design of numerous negative prompts to avoid generating low-quality results. Furthermore, the text embedding of ”low quality” might not be well disentangled during training, leading to an suboptimal results. Furthermore, while alternative guidance strategies (e.g. Autoguidance [Karras et al., 2024]) are capable of improving generation quality, they often come at the cost of significantly increased model parameters and sampling steps. In contrast, by incorporating trainable quality awareness, we are able to enhance overall generation quality without sacrificing sampling steps and inference memory consumption.